OpenStudy (anonymous):

use the squared identites to simplify: sin^2x cos^2x ?

OpenStudy (anonymous):

sin(2x)^2/4?

OpenStudy (anonymous):

A)\[1-\cos(4x)/4\] B)\[1+4\cos(2x)+\cos(4x)/4\] C)\[1+\cos(4x)/4\]

OpenStudy (anonymous):

a

OpenStudy (anonymous):

D)\[1+4\cos(2x)-\cos(4x)/4\]

OpenStudy (anonymous):

thanks :)